# GazeProphet: Software-Only Gaze Prediction for VR Foveated Rendering

### v1

@@ Abstract
Foveated rendering significantly reduces computational demands in virtual reality applications by concentrating rendering quality where users focus their gaze. Current approaches require expensive hardware-based eye tracking systems, limiting widespread adoption due to cost, calibration complexity, and hardware compatibility constraints. This paper presents GazeProphet, a software-only approach for predicting gaze locations in VR environments without requiring dedicated eye tracking hardware. The approach combines a Spherical Vision Transformer for processing 360-degree VR scenes with an LSTM-based temporal encoder that captures gaze sequence patterns. A multi-modal fusion network integrates spatial scene features with temporal gaze dynamics to predict future gaze locations with associated confidence estimates. Experimental evaluation on a comprehensive VR dataset demonstrates that GazeProphet achieves a median angular error of 3.83 degrees, outperforming traditional saliency-based baselines by 24% while providing reliable confidence calibration. The approach maintains consistent performance across different spatial regions and scene types, enabling practical deployment in VR systems without additional hardware requirements. Statistical analysis confirms the significance of improvements across all evaluation metrics. These results show that software-only gaze prediction can work for VR foveated rendering, making this performance boost more accessible to different VR platforms and apps.
Index Terms:
Virtual Reality, Gaze Prediction, Foveated Rendering, Vision Transformer, Eye Tracking
@@ I Introduction
The contributions of this work include: (1) a novel software-only approach for VR gaze prediction that eliminates hardware requirements, (2) integration of spherical vision transformers with temporal sequence modeling for VR environments, (3) comprehensive experimental validation showing practical performance levels, and (4) demonstration of reliable confidence estimation for adaptive foveated rendering applications.
@@ III-A Architecture Overview
The system outputs a 2D gaze prediction $(x,y$ coordinates)) and an associated confidence score. The confidence score indicates prediction reliability for adaptive foveated rendering applications.
@@ III-D3 Loss Function Design
Training optimizes a combined loss function balancing gaze accuracy and confidence calibration:
@@
Confidence loss ensures well-calibrated uncertainty estimates:
@@
where $c_{i}$ represents predicted confidence, $\tau$ sets accuracy threshold, and $\mathbb{I}$ is 1 if the condition is true and 0 otherwise.
Loss weights $\lambda_{gaze}=1.0$ and $\lambda_{conf}=0.1$ balance the two objectives during training.
The accuracy threshold $\tau=0.05$ corresponds to approximately 10 pixels on a 256×512 equirectangular image. This threshold balances prediction precision with practical VR requirements. Errors below this threshold produce imperceptible differences in foveated rendering quality.
Loss weight selection follows established multi-task learning principles. The gaze prediction weight $\lambda_{gaze}=1.0$ prioritizes spatial accuracy as the primary objective. The confidence weight $\lambda_{conf}=0.1$ provides sufficient regularization for uncertainty calibration without overwhelming the main prediction task. This 10:1 ratio prevents confidence learning from interfering with gaze accuracy optimization.
Preliminary experiments confirmed these values through validation performance analysis. Alternative weight ratios (1:1, 5:1) resulted in either degraded gaze accuracy or poorly calibrated confidence estimates.
@@ IV-A Performance Comparison
The confidence calibration demonstrates reliable uncertainty estimation. Average confidence scores of 0.997 correlate strongly with actual prediction accuracy. High confidence predictions (>0.9) achieve 97.1% accuracy, while lower confidence predictions show appropriately reduced performance.
@@ IV-D Statistical Significance
Statistical analysis confirms the significance of performance improvements across all metrics. Paired t-tests between GazeProphet and each baseline approach yield p-values below 0.001, indicating strong statistical significance.
@@ V Conclusion
This work presents GazeProphet, a software-only approach for gaze prediction in VR environments that enables foveated rendering without dedicated eye tracking hardware. The system combines a Spherical Vision Transformer for 360-degree scene analysis with an LSTM-based temporal encoder for gaze sequence modeling. Multi-modal fusion integrates these components to produce accurate gaze predictions with reliable confidence estimates.
@@
VI Future Work
Several research directions could further enhance VR gaze prediction capabilities. Head movement integration represents the most promising extension. VR environments naturally couple head and eye movements, providing additional predictive signals for gaze location estimation. Users typically turn their heads toward regions of interest before directing their gaze, creating predictive opportunities.
Real-time optimization remains essential for practical VR deployment. Model compression and specialized hardware acceleration could reduce latency below the 10 millisecond threshold required for seamless foveated rendering. User adaptation mechanisms could improve prediction accuracy through personalized gaze pattern learning over time.
Multi-user prediction in social VR environments and content-aware approaches for different VR application types represent additional research opportunities. These extensions could further improve prediction accuracy and expand the practical applications of software-only gaze prediction in VR systems.

### v2

@@ Abstract
Foveated rendering significantly reduces computational demands in virtual reality applications by concentrating rendering quality where users focus their gaze. Current approaches require expensive hardware-based eye tracking systems, limiting widespread adoption due to cost, calibration complexity, and hardware compatibility constraints. This paper presents GazeProphet, a software-only approach for predicting gaze locations in VR environments without requiring dedicated eye tracking hardware. The approach combines a Spherical Vision Transformer for processing 360-degree VR scenes with an LSTM-based temporal encoder that captures gaze sequence patterns. A multi-modal fusion network integrates spatial scene features with temporal gaze dynamics to predict future gaze locations. Experimental evaluation on a comprehensive VR dataset demonstrates that GazeProphet achieves a median angular error of 3.83 degrees, outperforming traditional saliency-based baselines by 24%. The approach maintains consistent performance across different spatial regions and scene types, enabling practical deployment in VR systems without additional hardware requirements. Statistical analysis confirms the significance of improvements across all evaluation metrics. These results show that software-only gaze prediction can work for VR foveated rendering, making this performance boost more accessible to different VR platforms and apps.
@@ I Introduction
The contributions of this work include: (1) a novel software-only approach for VR gaze prediction that eliminates hardware requirements, (2) integration of spherical vision transformers with temporal sequence modeling for VR environments, and (3) comprehensive experimental validation showing practical performance levels for foveated rendering applications.
@@ III-A Architecture Overview
The system outputs a 2D gaze prediction $(x,y$ coordinates)) and an associated confidence score for potential use in adaptive rendering strategies.
@@ III-D3 Loss Function Design
Training optimizes a combined loss function balancing gaze accuracy and confidence estimation:
@@
Confidence loss encourages prediction reliability awareness:
@@
where $c_{i}$ represents predicted confidence, $\tau$ sets the accuracy threshold, and $\mathbb{I}$ is 1 if the condition is true and 0 otherwise.
Loss weights $\lambda_{gaze}=1.0$ and $\lambda_{conf}=0.1$ balance the two objectives during training. The gaze prediction weight prioritizes spatial accuracy as the primary objective. The confidence weight provides sufficient regularization for uncertainty estimation without overwhelming the main prediction task. This 10:1 ratio prevents confidence learning from interfering with gaze accuracy optimization.
The accuracy threshold $\tau=0.05$ corresponds to approximately 10 pixels on a 256×512 equirectangular image. Preliminary experiments confirmed these values through validation performance analysis. Alternative weight ratios (1:1, 5:1) resulted in either degraded gaze accuracy or poorly calibrated confidence estimates.
@@ IV-A Performance Comparison
The acceptability of this 3.83 degree median angular error for practical foveated rendering depends on application requirements. Hardware-based eye tracking systems typically achieve sub-degree accuracy [4], though foveated rendering systems often use larger foveal regions (e.g., 15 degree radius) to account for tracking latency and display characteristics. Our software-only approach trades precision for accessibility, enabling foveated rendering on headsets without eye-tracking hardware. Whether this tradeoff is acceptable requires empirical validation in real-world VR applications.
@@ IV-D Statistical Significance
Statistical analysis confirms the significance of performance improvements across all metrics. Paired t-tests between GazeProphet and each baseline approach yield p-values below 0.001, indicating strong statistical significance. To account for multiple comparisons across metrics, we note that all reported p-values $(p<0.001)$ remain significant even under conservative Bonferroni correction ($\alpha=0.05/5=0.01$ for five primary metrics).
@@ V Conclusion
This work presents GazeProphet, a software-only approach for gaze prediction in VR environments that enables foveated rendering without dedicated eye tracking hardware. The system combines a Spherical Vision Transformer for 360-degree scene analysis with an LSTM-based temporal encoder for gaze sequence modeling. Multi-modal fusion integrates these components to produce accurate gaze predictions.
@@
VI Limitations
Inference latency. We did not measure inference latency. Real-time foveated rendering requires latency below 10 milliseconds, which needs validation on target VR hardware.
Single dataset evaluation. Evaluation relies on the Sitzmann VR Saliency dataset. Validation across additional VR datasets with different task types would strengthen generalizability claims.
Statistical analysis. Multiple comparison corrections were not applied to reported p-values. While our $p<0.001$ results remain significant under conservative Bonferroni correction, future work should explicitly apply these adjustments.
Practical thresholds. Whether 3.83 degrees median angular error suffices for practical foveated rendering depends on application requirements. Prior work assumes hardware eye tracking with sub-degree accuracy. Our software-only approach trades precision for accessibility. Empirical validation in real VR applications is needed to determine acceptable error ranges.
VII Future Work
Several research directions could enhance VR gaze prediction capabilities. Head movement integration represents a natural extension, as VR environments couple head and eye movements. Users often turn their heads toward regions of interest before directing gaze, potentially providing predictive signals. However, the benefit of head movement data for software-only gaze prediction requires empirical validation.
Real-time optimization remains essential for practical VR deployment. Model compression and hardware acceleration could reduce latency below the 10 millisecond threshold required for foveated rendering. Personalized adaptation mechanisms could improve accuracy through individual gaze pattern learning.
Validation across diverse VR applications (interactive gaming, social VR, dynamic content) and multiple datasets would establish generalizability. Improving confidence calibration for uncertainty-aware rendering decisions represents another important direction. These extensions could expand the practical applications of software-only gaze prediction in VR systems.